Documenta Mathematica

Band 18, 2013

Matteo Longo, Marc-Hubert Nicole The Λ-Adic

Shimura-Shintani-Waldspurger Correspondence  1–21

Hélène Esnault, Adrian Langer On a Positive Equicharacteristic

Variant of the p-Curvature Conjecture  23–50

Pablo Pelaez Birational Motivic Homotopy Theories

and the Slice Filtration  51–70

André Joyal and Joachim Kock

Coherence for Weak Units  71–110

iv